(a) list the fundamental quantitis and their units .
(b) A natural gas has following composition by volume . methane=80% ;ethane=12% ;butane =3% and the rest nitrogen gas .
(1) calculate the composition of the gas by mass.(2) calculate its average molecular weight .(3) calculat the density of the gas at NTP
1
Expert's answer
2020-05-04T01:53:25-0400

a) Quantity is s property that is measured [e.g. mass, length, time, volume, pressure]. Unit is a standard quantity against which a quantity is measured [e.g. gram, metre, second, litre, pascal; which are units of the above quantities]. Chemists measure various quantities. 

b) 1) If we have 1L of gas:

n(CH4) = V/Vm = 0.8L/22.4L/mol = 0.0357 mol

m(CH4) = n*M = 0.0356mol*16g/mol = 0.5712 g

n(C2H6) = 0.12L/22.4L/mol = 0.00536 mol

m(C2H6) =0.00536mol*30g/mol = 0.161 g

n(C4H10) = 0.03L/22.4L/mol = 0.00134 mol

m(C4H10) = 0.00134mol*58g/mol = 0.078 g

n(N2) = 0.05L/22.4L/mol = 0.00223 mol

m(N2) = 0.00223mol*28g/mol = 0.062 g

2) Average molecular weight = 16*0.8+30*0.12+58*0.03+28*0.05 = 19.54 g/mol

3) Density of the gas = Mav/Vm = 19.54g/mol/22.4L/mol = 0.872 g/L